OpenStudy (anonymous):

Solve the logarithmic equation. log10(x) = –5/6

OpenStudy (anonymous):

is that base 10?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

is this \[log(10^x)=-\frac{5}{6}\]

OpenStudy (anonymous):

or \[log_{10}(x)=-\frac{5}{6}\]?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

which one? i wrote two different ones

OpenStudy (anonymous):

last one

OpenStudy (anonymous):

then is says \[x=10^{-\frac{5}{6}}\]

OpenStudy (anonymous):

you haven't soldit

OpenStudy (anonymous):

you haven't soldit

OpenStudy (anonymous):

you haven't soldit

OpenStudy (anonymous):

you haven't soldit

OpenStudy (anonymous):

because i don't have a calculator handy. i have no idea what \[10^{-\frac{5}{6}}\] is